OpenStudy (anonymous):

what is tan theta = 1

OpenStudy (anonymous):

theta is the angle. hence tan theta= 1 implies theta=45 degrees.

OpenStudy (anonymous):

how do you know that thought? is there a formula to get it?

OpenStudy (anonymous):

\[45degress = \pi/4\]

OpenStudy (anonymous):

so the answer is 45?

OpenStudy (anonymous):

it depends if the question is asking for degrees or radians.

OpenStudy (anonymous):

it asks for radians. so how did you get from tan (theta) = 1 to 45? thats what im confused on, where does the 45 come from?

OpenStudy (anonymous):

45 degrees is the angel at which makes \[\tan \theta =1\] so \[\tan 45^{0} = 1\]

OpenStudy (anonymous):

\[\theta=\tan^{-1}(1) \]

OpenStudy (anonymous):

ohhhh okay! thanks so much! so if you calculate tan^(-1) you get 45! okay thanks so much :)

OpenStudy (anonymous):

\[\theta=45 \deg or \pi/4 radians\]

OpenStudy (anonymous):

one more quick thing! is there a possibility for a second answer??

OpenStudy (anonymous):

hey there is conversion from degrees to radians. \[\pi=180 \deg\]

OpenStudy (anonymous):

\[\pi radians = 180 \deg\]

OpenStudy (anonymous):

so you can simply convert from radians to degrees or degrees to radian.

OpenStudy (anonymous):

okay thank you! also i see that in the answer booklet i have that there are two answers to this question pi/4 and 5pi/4 . where does the 5pi/4 come from??

OpenStudy (anonymous):

both are only the different representation schemes. only matter of unit radian or degree.

OpenStudy (anonymous):

okay, thank you!

OpenStudy (anonymous):

it will be clear only when you will type the problem. but what i see from the answer is that the signal you are dealing with is a periodic signal.

OpenStudy (anonymous):

by i am leaving .... All the best

OpenStudy (anonymous):

thank you:)